OpenStudy (anonymous):

The width of a rectangle is (3x − 1) cm. The length of the rectangle is twice the width. Find the area of the rectangle in terms of the variable x.

OpenStudy (lilg132):

2(3x-1) = (6x - 2) (6x - 2) (3x - 1) = Area

OpenStudy (anonymous):

The area of a rectangle is the product of the width and the length. You already have the width (3x - 1). You know the length is twice this. How would you write the length?

OpenStudy (anonymous):

2(3x-1)^2

OpenStudy (anonymous):

That's a perfect answer as far as I'm concerned. Well done!

OpenStudy (anonymous):

\[2(3x-1)(3x-1) = 2(3x-1)^2\]The answers are equally valid without some kind of condition in the question.

OpenStudy (lilg132):

wouldnt that mean (6x - 2)^2 ? the same as (6x - 2) (6x - 2)

OpenStudy (anonymous):

Conventionally, the square would only apply to the bracket, not the two. If you wanted the 2 to be squared, you would normally write: \[2^2(3x-1)^2\]or\[(2(3x-1))^2\]

OpenStudy (lilg132):

ok i get itt thanks